# GlossLM: Multilingual Pretraining for Low-Resource Interlinear Glossing

### v1

@@ 6 Results
Figure 2: Word accuracy for the multilingual GlossLM and various baseline systems on unsegmented data for out-of-domain languages.
Accuracy results for the segmented setting are given in Table 3, and for the unsegmented setting in Table 3 and Figure 2. We report chrF++ scores in Appendix D, finding that they generally show the same trends as accuracy. [^5] All results are for a single run, due to the computational cost of multiple runs.
Segmented setting
In the segmented setting, we observe mixed results. The GlossLM models outperform baselines for certain languages (arp, usp, git) but underperform in other cases (ddo, lez, ntu, nyb). GlossLM matches or outperforms the comparable ByT5 baseline, indicating that the continual gloss pretraining provides small benefits. However, this method does not necessarily overcome other architectures that are more successful, such as the straight-through gradient model Girrbach (2023a) (though future work could utilize our corpus to pretrain models with these architectures). We note that for this setting, performance across models is very high, and it may be difficult to achieve much higher than the observed performances (due to ambiguous morphemes or noisy evaluation data).
Notably, the direct ByT5 model trained on Gitksan far underfits the data, producing improperly formatted data, long random sequences, and outputs including chunks of the prompts. As the Gitksan dataset is very small, the model was unable to learn the correct formatting; meanwhile, the GlossLM model benefits from learning correct formatting from other languages, for a 26.7% morpheme accuracy improvement. For very low-resource documentation projects, our pretrained model shows clear benefits in producing outputs with the correct formatting, while leveraging the multilingual knowledge of ByT5.
We also observe that while finetuning on in-domain languages provides small improvements in performance, the non-finetuned models perform very well. This indicates that multilinguality has not limited performance on each individual language too much, despite the hugely multilingual pretraining corpus. Furthermore, the non-finetuned model is able to make a small percentage of correct predictions (2.9% to 4.9% morpheme accuracy) on unseen languages without training, indicating clear benefits from massively multilingual pretraining.
Unsegmented setting
In the unsegmented setting, the GlossLM models outperform most prior baselines on the in-domain data, although they slightly underperform the ByT5 baseline. This result indicates that with large monolingual datasets, there is not necessarily a benefit to additional multilingual learning on IGT data. This could indicate that the multilingual (non-IGT) corpus used to pretrain ByT5 already sufficiently provides any benefits from crosslingual transfer. The non-finetuned model is also highly competitive, beating prior baselines (except the ByT5 baseline).
For the out-of-domain languages (which tend to have much smaller corpora), the finetuned GlossLM model outperforms the ByT5 baseline on three of four languages and beats previous best baselines by up to 6.6% morpheme accuracy (for Natugu). When the monolingual dataset is very limited, there seems to be a greater benefit from multilingual IGT pretraining, allowing the model to leverage information about segmentation and output formatting from languages in the pretraining corpus. The only case where the finetuned model underperforms SOTA is for Lezgi, slightly underperforming the Tü-CL model. One possibility is that segmentation from surface forms in Lezgi is more difficult, and the hard attention mechanism is beneficial at learning proper morpheme boundaries.
Overall, we suggest the pretrained GlossLM model is a good fit for languages with limited training data, even unseen languages. Directly finetuning a ByT5 model can be very effective for languages with large training corpora, but few documentation projects will have as much available data as the in-domain languages used here (which have 7k to 78k training sentences), and finetuning directly on small datasets risks failure to converge.
Furthermore, we observe that finetuning the GlossLM model on unseen data converges much faster than ByT5–for example, after 50 train epochs for Natugu, the GlossLM-based model achieves 78.1 chrF++ on the eval set, while the ByT5 model is at 30.1. If the desire is to conserve computational resources, using the pretrained model could be beneficial to train efficiently.
7 Error Analysis
7.1 In- vs out-of-vocabulary errors
In-domain Out-of-domain
arp ddo usp git lez ntu nyb
% OOV 30.0 15.6 20.0 78.1 27.3 27.6 8.42
IV 96.2 92.3 91.4 66.7 85.4 91.3 92.8
OOV 55.7 71.7 57.1 26.0 33.9 55.7 32.6
% OOV 30.0 18.7 21.4 80.5 25.5 28.9 9.27
IV 95.3 91.4 89.1 60.0 81.6 91.3 91.7
OOV 50.1 69.7 50.9 20.7 16.4 48.3 30.6
We are interested in how well our models can generalize to unseen words, which are often novel combinations of new and previously-seen morphemes. We report word-level accuracy for our finetuned GlossLM models, indicating whether the transcribed word is in- or out-of-vocabulary in Table 4, as well as the percent of OOV words in the test set. The OOV rate between segmented and unsegmented may vary slightly, as mappings between segmented and unsegmented forms are not necessarily one-to-one. We consider a word to be in-vocabulary if the form of the word in the transcription and its corresponding gold label in the gloss co-occur in the training data. We also include morpheme-OOV rates and statistics on lexical gloss overlap with translations in Appendix E.
Across most languages, our models perform well on in-vocabulary words (with Gitksan being a notable exception, likely due to the small amount of training data). Within the OOV split, we expect that models perform better on languages that are agglutinative with lower morpheme-OOV rates. We find that this is the case, as Tsez (which has the highest OOV accuracy) is a highly agglutinative language with the lowest rate of OOV morphemes in the test set, implying that the majority of OOV words are actually composed of seen morphemes. However, Lezgi—which is also highly agglutinative—has strikingly low OOV accuracy; this may be due to the highly complex inflections for case marking, which could be more difficult for our system to learn. Finally, Nyangbo also shows relatively low OOV word accuracy, possibly due to the lack of translations in the input, which could help predict OOV stem labels.
7.2 Qualitative Analysis of Errors
A manual analysis of GlossLM outputs reveal two interesting points of consideration.
In certain cases, the model may rely too heavily on translations when predicting lexical glosses.
When inspecting the Uspanteko outputs from the finetuned GlossLMall model, we found that the model outputs “esta@bien” (the correct label for the morpheme tziin) as the label for jaan, which should be glossed as “bueno”; this happens especially often when the phrase “esta bien” appears in the translation. In other cases, the model sometimes outputs a word or phrase present in the translation, even if there is no occurrence of that item in the gold gloss. We include example system outputs to illustrate these cases in Appendix F.
Our metrics potentially underestimate GlossLM’s performance on the glossing task.
There are sometimes inconsistencies in the IGT labels where multiple interchangeable glosses are used for the same morpheme, particularly for lexical glosses (e.g. the Arapaho word ’eeneisih’i is glossed in the data as “how.X.things.are.named”, “how.s.o..is.named”, and “how.things.are.named”). These types of errors can be accounted for when using chrF++, but another issue that may more greatly impact generative models—especially if it learns to use translations effectively—is if the model outputs a semantically similar label that is dissimilar in form; for example, outputting “PREP” (preposition label) instead of “en” (“in” in Spanish, which is a preposition). We include examples of such outputs in Appendix F.
@@ 8 Related Works
8.1 Automatic Interlinear Glossing
@@
8.2 Large Multilingual Pretrained Models
@@ 9 Conclusion
Automatic generation of interlinear glossed text has great potential to aid documentation projects, but limited monolingual data limits effectiveness. In this research, we attempt to overcome these limitations using crosslingual transfer, compiling the largest multilingual IGT dataset and pretraining massively multilingual glossing models. Our models outperform any existing work in the majority of settings, demonstrating clear benefits of multilingual representations–particularly for languages with very small corpora. We make our dataset and models available for future research, in hopes that they will aid future NLP research and benefit documentation projects.
10 Limitations
@@ Appendix C Training Hyperparameters
Appendix D chrF++ Scores
We provide chrF++ scores for various systems in Table 10. Broadly, we found the results supported the same trends as the accuracy measurements, although the magnitude of the differences varied.
Appendix E Test Split Vocabulary Statistics
@@
We include additional statistics for morpheme OOV rate and recall rate for lexical labels in the translation in Table 7 and Table 8, as reported in Ginn et al. (2023).
@@ Appendix F Example System Outputs
We include examples of system outputs to illustrate the types of errors discussion in subsection 7.2 in Table 9.
Model chrF++ Score (Segmented)
In-domain Out-of-domain
arp ddo usp Avg git lez ntu nyb Avg
Top-choice 75.0 71.9 71.4 72.8 33.7 69.7 74.5 62.2 60.0
Token-class 84.2 81.8 75.3 80.4 25.6 52.1 65.4 84.3 56.9
Tü-CL 85.2 88.4 77.7 83.8 34.8 78.9 87.9 87.7 72.3
CRF 84.2 88.2 79.4 83.9 40.9 79.2 88.4 84.3 73.2
SMF 80.7 86.6 76.3 81.2 28.0 62.0 78.8 82.2 62.6
ByT5all 84.2 91.8 84.6 86.9 9.07 74.9 85.1 88.8 64.5
GlossLMall, pre 85.2 90.8 83.1 86.4 21.5 14.1 18.2 8.8 15.7
GlossLMall, ft 86.3 91.5 85.9 87.9 43.1 75.0 84.1 87.4 72.4
Model chrF++ Score (Unsegmented)
In-domain Out-of-domain
arp ddo usp Avg git lez ntu nyb Avg
Top-choice 44.0 63.5 55.1 54.2 8.4 51.6 40.5 74.0 43.6
Token-class 56.2 72.9 65.3 64.8 18.8 56.4 45.1 18.8 34.8
Tü-CL 77.6 84.6 72.5 78.2 23.0 71.5 78.6 84.1 64.3
ByT5unseg 80.7 90.0 83.0 84.6 7.6 59.6 77.0 88.4 58.2
GlossLMunseg, pre 80.5 86.8 81.0 82.8 19.4 13.3 16.3 8.1 14.3
GlossLMunseg, ft 82.9 89.8 81.7 84.8 34.9 68.8 80.7 85.5 67.5

### v2

@@ 7 Results
7.1 Comparison with Baselines
Figure 4: Morpheme accuracy for various systems.
After pretraining GlossLM, we train finetuned versions for each of the languages in the test set. We first describe our finetuning procedure and compare results of our finetuned models against baselines from Ginn (2023)7.1). Then, to further isolate the efficacy of pretraining, we compare the finetuned versions of GlossLM to a finetuned ByT5 model without the multilingual gloss pretraining (§7.2). Finally, we explore whether training on a minimally normalized version of the data improves performance (§7.3).
As previously noted, we focus on the unsegmented setting; for completeness, we provide full results for both segmented and unsegmented data in Appendix D.
Finetuning can help align the model to a particular language or even a new unseen language. We finetune our GlossLM pretrained model on the training dataset for each language individually, and label these runs as GlossLMft. We do this for both the in-domain languages, to focus the model on a single language, as well as the out-of-domain languages, allowing us to study the model’s adaptation to unseen languages.
Finetuning used the same parameters as pretraining, but with 100 training epochs and early stopping (patience 3, start epoch 15) [^5] For Gitksan, due to the size of the training set, we set max epochs to 300 and patience to 15., and took anywhere from 20 minutes to one day for each language. Inference uses beam search with $n=3$ beams.
We compare the finetuned GlossLM models with three baseline systems which include the state-of-the-art from prior work:
  1. 1.
    Top-choice selects the most frequent label associated with each morpheme/word in the training data, and assigns “???” to unseen morphemes.
  2. 2.
    Token-class treats the glossing task as a token classification problem, where the output vocabulary consists of the IGT morpheme or word-level labels. Each target language’s data is used to train a language-specifc Token-class model, which uses the RoBERTa architecture with default hyperparameters without any additional pretraining Liu et al. (2019). This was used as the baseline model for the SIGMORPHON 2023 Shared Task on interlinear glossing Ginn (2023).
  3. 3.
    Tü-CL Girrbach (2023b) uses straight-through gradient estimation (Bengio et al., 2013) to induce latent, discrete segmentations of input texts, and predicts glosses using a multilayer perceptron.
As shown in  Figure 4, we find that our finetuned models outperform SOTA in all but two languages (Gitksan and Lezgi). The Tü-CL model (Girrbach, 2023a), which is a close second and outperforms on Gitksan and Lezgi, uses explicit latent segmentations, which seems to be particularly beneficial for the very low-resource, unsegmented setting.
7.2 Comparison with Finetuned ByT5
To directly assess the impact of pretraining on performance, we finetune ByT5 models on each language in the test set with the same configuration as for the finetuned GlossLM models. We then compare the performance of our models (which have under gone both multilingual gloss pretraining and finetuning) with analogous ByT5 models (without multilingual gloss pretraining), as shown in Figure 5.
Figure 5: Performance after monolingual finetuning, comparing a standard pretrained ByT5 with a continually pretrained GlossLM model. The x-axis uses the log (base 10) of the number of training examples in a given language, for readability.
We observe mixed results, which are largely dependent on the size of the training corpus. For languages with less training data (Gitksan, Lezgi, Natugu) the GlossLMft model outperforms the finetuned ByT5 model (by 10.0, 15.1, and 9.1 points respectively). In the case of Gitksan, the finetuned ByT5 model is completely unable to produce well-formatted output (likely due to the tiny training corpus) while the GlossLMft model does not struggle with this as much. A possible explanation is that even if there are no similar languages in the pretraining corpus, the GlossLMft can leverage knowledge about IGT formatting from unrelated languages.
With enough data (starting with Nyangbo, 2100 examples), the two approaches achieve nearly identical performance. This is an unsurprising result, indicating that large amounts of high-quality monolingual data overshadow any benefits from crosslingual transfer. Furthermore, we note that benefits of multilingual gloss pretraining shown may not be unique to the T5 architecture—while we only experiment with ByT5, our pretraining strategy could be applied to other architectures.
7.3 Effect of Gloss Normalization
Finally, we experiment with pretraining and finetuning on a minimally normalized version of the dataset, where the 200 most frequent grammatical labels are mapped to a set of standard labels.
We repeat the same pretraining and finetuning process as before. When comparing the performance of the pretrained model before finetuning on in-domain languages, we find minimal differences compared to the results in §6.3. This is also the case after finetuning, though we found improvements in three of the four out-of-domain languages: Lezgi ($+\Delta 2.9$), Natugu ($+\Delta 6.1$), and Nyangbo ($+\Delta 2.5$), which (aside from Gitksan) are the three lowest resourced languages in our test set, exceeding current SOTA performance. From this, it appears that normalization may be most helpful when adapting the pretrained model to an unseen language with a low-to-moderate amount of training data.
@@ 8 Related Works
Automatic Interlinear Glossing
@@
Large Multilingual Pretrained Models
@@ 9 Conclusion
High-quality language documentation involves an incredible amount of effort. We compile, normalize, and release the largest corpus of multilingual IGT data, enabling future research in linguistics, NLP, and documentation. Furthermore, we demonstrate the applicability of our corpus by pretraining a multilingual neural model for automatic generation of IGT. We finetune the model on monolingual corpora, showing benefits on low-resource languages due to multilingual pretraining. In five out of seven languages, we achieve a new SOTA on automatic IGT generation.
Limitations
@@ Appendix C Training Hyperparameters
Appendix D Full Results
We provide full results for accuracy and chrF++ scores in Table 5 and Table 6 . For our normalization experiments, we only trained and tested on unsegmented data for the target languages.
Model Morpheme accuracy / Word accuracy (Segmented)
In-domain Out-of-domain
arp ddo usp Avg git lez ntu nyb Avg
Model Morpheme accuracy / Word accuracy (Unsegmented)
In-domain Out-of-domain
arp ddo usp Avg git lez ntu nyb Avg
GlossLMunseg, ft 82.1 / 81.5 83.6 / 87.3 78.6 / 81.0 81.4 / 83.3 10.1 / 28.4 57.3 / 64.9 62.8 / 78.9 87.4 / 86.2 54.4 / 64.6
GlossLM-normunseg, pre 79.6 / 80.0 79.6 / 83.2 74.8 / 76.6 78.0 / 79.9 2.2 / 7.8 2.6 / 1.8 2.9 / 9.7 1.0 / 2.5 2.2 / 5.45
GlossLM-normunseg, ft 82.0 / 81.5 84.2 / 87.8 76.4 / 79.2 80.8 / 82.8 9.3 / 16.4 60.3 / 67.8 63.4 / 76.6 90.0 / 87.6 55.8 / 62.1
Model chrF++ Score (Segmented)
In-domain Out-of-domain
arp ddo usp Avg git lez ntu nyb Avg
Top-choice 75.0 71.9 71.4 72.8 33.7 69.7 74.5 62.2 60.0
Token-class 84.2 81.8 75.3 80.4 25.6 52.1 65.4 84.3 56.9
Tü-CL 85.2 88.4 77.7 83.8 34.8 78.9 87.9 87.7 72.3
CRF 84.2 88.2 79.4 83.9 40.9 79.2 88.4 84.3 73.2
SMF 80.7 86.6 76.3 81.2 28.0 62.0 78.8 82.2 62.6
ByT5all 84.2 91.8 84.6 86.9 9.07 74.9 85.1 88.8 64.5
GlossLMall, pre 85.2 90.8 83.1 86.4 21.5 14.1 18.2 8.8 15.7
GlossLMall, ft 86.3 91.5 85.9 87.9 43.1 75.0 84.1 87.4 72.4
Model chrF++ Score (Unsegmented)
In-domain Out-of-domain
arp ddo usp Avg git lez ntu nyb Avg
Top-choice 44.0 63.5 55.1 54.2 8.4 51.6 40.5 74.0 43.6
Token-class 56.2 72.9 65.3 64.8 18.8 56.4 45.1 18.8 34.8
Tü-CL 77.6 84.6 72.5 78.2 23.0 71.5 78.6 84.1 64.3
ByT5unseg 80.7 90.0 83.0 84.6 7.6 59.6 77.0 88.4 58.2
GlossLMunseg, pre 80.5 86.8 81.0 82.8 19.4 13.3 16.3 8.1 14.3
GlossLMunseg, ft 82.9 89.8 81.7 84.8 34.9 68.8 80.7 85.5 67.5
GlossLM-normunseg, pre 80.3 86.6 78.7 81.9 19.5 14.3 16.6 7.4 14.5
GlossLM-normunseg, ft 82.6 89.9 81.1 84.5 29.1 70.9 79.0 87.5 66.6
Appendix E In- vs out-of-vocabulary errors
In-domain Out-of-domain
arp ddo usp git lez ntu nyb
% OOV 30.0 15.6 20.0 78.1 27.3 27.6 8.42
IV 96.2 92.3 91.4 66.7 85.4 91.3 92.8
OOV 55.7 71.7 57.1 26.0 33.9 55.7 32.6
% OOV 30.0 18.7 21.4 80.5 25.5 28.9 9.27
IV 95.3 91.4 89.1 60.0 81.6 91.3 91.7
OOV 50.1 69.7 50.9 20.7 16.4 48.3 30.6
We report word-level accuracy for our finetuned GlossLM models, indicating whether the transcribed word is in- or out-of-vocabulary in Table 7, as well as the percent of OOV words in the test set. The OOV rate between segmented and unsegmented may vary slightly, as mappings between segmented and unsegmented forms are not necessarily one-to-one. We consider a word to be in-vocabulary if the form of the word in the transcription and its corresponding gold label in the gloss co-occur in the training data. We also include morpheme-OOV rates and statistics on lexical gloss overlap with translations in in Table 8 and Table 9, as reported in Ginn et al. (2023).
@@ Appendix F Example System Outputs
We include examples of system outputs to illustrate the types of errors discussion in Table 10.